A private investment club has $175,000 e been classified into three categories: high-risk, medium-risk, and low-risk. Management estimates t stocks, 10%/year, and low-risk stocks, 6%/year. The members have of the other two categories. Determine how much the club should invest in each type of stock if the investment goal is to have a return of $17,500/year on the total for investment in stocks. To arrive at an acceptable overall level of risk, the stocks that management is considering have hat high-risk stocks will have a return rate of 15%/year medium-risk decided that the investment in low-risk stocks should be equal to the sum of the investments in the stocks available for investment is invested. Let x, y, and z denote the amount, in dollars, invested in high-, medium, and low-risk stocks, 175,000 - 17,500
Answer #1

Please provide rating...

Equation formation
x+y+z = 175000
x+y= z
x*15%+y*10%+z*6%= 17500